OpenStudy (anonymous):

What is the slant height if the surface area is 395.64 square meters? (drawing)

OpenStudy (anonymous):



OpenStudy (anonymous):

is it total surface area of curved surface area???

OpenStudy (anonymous):

It doesn't say. All it says is (use 3.14 for n)

OpenStudy (anonymous):

If it is Curved Surface area then it is = πrl where l is the slant height Let the slant height be = l m ATQ, 395.64 sqm = 3.14 * 7 m * l m 395.64 ------- = l 3.14 * 7 hence l = 18 so the slant height is 18 m

OpenStudy (anonymous):

o: wait i think they mean the total surface area

OpenStudy (anonymous):

then total surface area = πrl + πr^2 = πr(l + r) 395.64 = 3.14*7 (l + 7) 395.64 ------- = l + 7 3.14*7 18 = l + 7 18 - 7 = l 11 = l so the slant height is 11 m is it ok now????

OpenStudy (anonymous):

Yess :D ! Ty

OpenStudy (anonymous):

u r welcome...☺